OpenStudy (anonymous):

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (g o f)(3)?

OpenStudy (anonymous):

(g o f)(3)= g(f(3) so first find f(3) then use the result in f(3) in g(x)

OpenStudy (anonymous):

hope u will be able to complete it...

OpenStudy (anonymous):

no I wont D: @matricked im completely dumb when it comes to this

OpenStudy (anonymous):

f(3) = -2*3+8=-6+8=2 now we need g(2) =3(2)=6

OpenStudy (anonymous):

See if you don't learn then simply knowing the answers won't help u out every time.. You can start learning as soon as possible so that next when u see numbers u don't be dumb..

OpenStudy (anonymous):

so what do we do after?

OpenStudy (anonymous):

f(3) = -2*3+8=-6+8=2 now we need g(2) =3(2)=6

OpenStudy (anonymous):

ok you have f and g so ????? what no @matricked

OpenStudy (anonymous):

I got 23???? @matricked